# Kernel Banzhaf: A Fast and Robust Estimator for Banzhaf Values

### v1

@@ 3.2 The Kernel Banzhaf Algorithm
Since $\mathbf{A}$ and $\mathbf{b}$ are exponentially large, constructing the linear regression problem to calculate the Banzhaf values is computationally prohibitive. Instead, we construct a smaller linear regression problem with $m$ samples. Let the subsampled matrix be $\tilde{\mathbf{A}}\in\mathbb{R}^{m\times n}$ and the target vector be $\tilde{\mathbf{b}}\in\mathbb{R}^{m}$. The estimate we produce is
$\hat{\bm{\phi}}=\operatorname*{arg\,min}_{\mathbf{x}}\|\tilde{\mathbf{A}}\mathbf{x}-\tilde{\mathbf{b}}\|_{2}.$ (8)
Building on this framework, we introduce the Kernel Banzhaf algorithm, which utilizes paired sampling and leverage score sampling to construct the small regression problem (Woodruff et al., 2014; Drineas & Mahoney, 2018). The leverage scores of $\mathbf{A}$ are given by
$\ell_{\mathbf{z}}=[\mathbf{A}]_{\mathbf{z}}\left(\mathbf{A^{\top}A}\right)^{-1}[\mathbf{A}]_{\mathbf{z}}^{\top}=\frac{1}{2^{n-2}}\left(\frac{1}{2}\mathbf{z}^{\top}\right)\left(\frac{1}{2}\mathbf{z}\right)=\frac{n}{2^{n}}$ (9)
where the third equality follows by Observation 3.1. Since the leverage scores are the same for all rows, uniform sampling is equivalent to sampling by leverage scores for $\mathbf{A}$. As such, Kernel Banzhaf selects $\tilde{\mathbf{A}}$ and $\tilde{\mathbf{b}}$ by sampling each row uniformly at random with replacement.
The pseudocode for Kernel Banzhaf appears in Algorithm 1.
@@
1:Set function $v:\{0,1\}^{n}\to\mathbb{R}$, $n$ players, $m$ samples s.t. $n<m\leq 2^{n}$.
2:Approximate Banzhaf values $\hat{\bm{\phi}}\in\mathbb{R}^{n}$
3:$\mathbf{Z}\leftarrow\mathbf{0}_{m\times n}$
4:for  $i=0,\ldots,\lfloor m/2\rfloor-1$  do
5:     $\mathbf{z}\leftarrow$ random binary vector sampled uniformly
6:     $\bar{\mathbf{z}}\leftarrow$ complement of $\mathbf{z}$triangleright I.e., $\bar{z}_{j}=1-z_{j}$ for $0\in\{0,\ldots,n-1\}$
7:     $\mathbf{Z}_{2i}\leftarrow\mathbf{z}^{\top}$, $\mathbf{Z}_{2i+1}\leftarrow\bar{\mathbf{z}}^{\top}$
8:end for
9:$\tilde{\mathbf{b}}\leftarrow[v(\mathbf{Z}_{1}),\ldots,v(\mathbf{Z}_{m})]$
10:$\tilde{\mathbf{A}}\leftarrow\mathbf{Z}-\frac{1}{2}\mathbf{1}_{m\times n}$triangleright Convert from $\{0,1\}$ to $\{-\frac{1}{2},\frac{1}{2}\}$
11:$\hat{\bm{\phi}}\leftarrow\operatorname*{arg\,min}_{\mathbf{x}}\|\tilde{\mathbf{A}}\mathbf{x}-\tilde{\mathbf{b}}\|_{2}$triangleright Standard least squares
12:return $\hat{\bm{\phi}}$
We now consider the time complexity of Algorithm 1. The time to randomly sample with replacement is about $O(n)$ per sample, along with the time to compute the complement and store both vectors. Let $T_{m}$ be the time complexity to evaluate the set function $m$ times, possibly in parallel. This cost is unavoidable for any algorithm that evaluates the set function $m$ times. Finally, the least squares regression problem requires $O(mn^{2})$ time because the matrix $\tilde{\mathbf{A}}$ has $m$ rows and $n$ columns and the vector $\tilde{\mathbf{b}}$ has $m$ entries. In total, the time complexity of Algorithm 1 is $O(T_{m}+mn^{2})$. For most set functions, we expect the $O(T_{m})$ term to dominate. For example, even a forward pass on a shallow fully connected neural network with dimension at most $n$ will take $O(n^{2})$ time per sample.
The state-of-the-art methods for Shapley estimation, KernelSHAP (Lundberg & Lee, 2017; Covert & Lee, 2020) and Leverage SHAP (Musco & Witter, 2024), sample without replacement to enhance performance when $m\approx n$. Our empirical results in Figure 10 show no improvements with this approach (until $m=2^{n}$, in which case we can exactly compute Banzhaf values). Moreover, sampling without replacement exhibits lower efficiency and scalability, and it does not ensure the use of exactly $m$ samples. Further details are provided in Appendix G.
@@ 3.3 Approximation Guarantees
If $m=O(n\log\frac{n}{\delta}+\frac{n}{\delta\epsilon})$, Algorithm 1 produces an estimate $\hat{\bm{\phi}}$ that satisfies
$\|\mathbf{A}\hat{\bm{\phi}}-\mathbf{A}\bm{\phi}\|_{2}\leq(1+\epsilon)\|\mathbf{A}\bm{\phi}-\mathbf{b}\|_{2}$ (10)

### v2

@@ 3.2 The Kernel Banzhaf Algorithm
Since $\mathbf{A}$ and $\mathbf{b}$ have $2^{n}$ rows, constructing the linear regression problem in Theorem 3.1 to calculate the Banzhaf values is computationally prohibitive. In particular, we must evaluate $v(S)$ for all $2^{n}$ subsets of $[n]$ to compute $\mathbf{b}$. Inspired by Kernel SHAP, we avoid this cost by constructing a much smaller regression problem that contains a subsample of $m\ll 2^{n}$ rows from $\mathbf{A}$ and $\mathbf{b}$. Let the subsampled design matrix be $\tilde{\mathbf{A}}\in\mathbb{R}^{m\times n}$ and the subsampled target vector be $\tilde{\mathbf{b}}\in\mathbb{R}^{m}$. The estimate we produce is
$\tilde{\boldsymbol{\phi}}=\operatorname*{arg\,min}_{\mathbf{x}}\|\tilde{\mathbf{A}}\mathbf{x}-\tilde{\mathbf{b}}\|_{2}^{2}.$
The most direct approach to constructing $\tilde{\mathbf{A}}$ and $\tilde{\mathbf{b}}$ would be to subsample $m$ rows from $\mathbf{A}$ and $\mathbf{b}$ uniformly at random. We do so with a slight modification – whenever the row corresponding to a set $S$ is sampled, we also select the row corresponding to $[n]\setminus S$. Referred to as “paired sampling”, this approach has been used in the Kernel SHAP algorithm for Shapley values (Covert & Lee, 2020). While it has no impact on our theoretical results (unpaired sampling would give almost identical bounds) we observe an experimental improvement for Banzhaf value estimation as well. See e.g, Figure 2, where unpaired sampling is denoted by “Kernel Banzhaf (excl. Pairs).” Detailed pseudocode for our Kernel Banzhaf appears in Algorithm 1. In Appendix G, we also explore the possibility of uniform sampling without replacement; ultimately, we find that sampling with replacement is simpler and offers similar performance, until $m\geq 2^{n}$ when we can exactly compute Banzhaf values anyways.
@@
  Input: $n$: positive integer, $v:\{0,1\}^{n}\to\mathbb{R}$: set function, $m$ : even number of samples such that $n<m\leq 2^{n}$
  Output: $\tilde{\boldsymbol{\phi}}\in\mathbb{R}^{n}$: estimated Banzhaf values
  Initialize $\tilde{\mathbf{A}}\leftarrow\mathbf{0}_{m\times n}$
  for $j\in\{0,2,4,\ldots,m\}$ do
     Sample $S_{j}$ uniformly from all subsets of $[n]$
     Set $\tilde{\mathbf{A}}_{j}\leftarrow\mathbbm{1}[S_{j}]-\frac{1}{2}\mathbf{1}$
     Compute $S_{j+1}=[n]\setminus S$ // Paired sampling
     Set $\tilde{\mathbf{A}}_{j+1}\leftarrow\mathbbm{1}[S_{j+1}]-\frac{1}{2}\mathbf{1}$
  end for
  Compute $\tilde{\mathbf{b}}\leftarrow[v(S_{1}),\ldots,v(S_{m})]$
  Solve $\tilde{\boldsymbol{\phi}}\leftarrow\operatorname*{arg\,min}_{\mathbf{x}}\|\tilde{\mathbf{A}}\mathbf{x}-\tilde{\mathbf{b}}\|_{2}$
  Return: $\tilde{\boldsymbol{\phi}}$
MC MSR Kernel Banzhaf
Dataset 25% Median 75% 25% Median 75% 25% Median 75%
Diabetes ($n=8$) 0.0053 0.0173 0.0357 0.0275 0.0368 0.0537 0.0002 0.0006 0.0011
Adult ($n=14$) 0.0075 0.0116 0.0263 0.0387 0.0482 0.0633 0.0004 0.0006 0.0010
Bank ($n=16$) 0.0106 0.0169 0.0448 0.0412 0.0512 0.0663 0.0009 0.0012 0.0022
German Credit ($n=20$) 0.0085 0.0158 0.0352 0.0399 0.0511 0.0682 0.0006 0.0009 0.0015
NHANES ($n=79$) 0.0008 0.0019 0.0038 0.0444 0.0496 0.0533 0.0000 0.0000 0.0000
BRCA ($n=100$) 0.0057 0.0114 0.0415 0.0479 0.0533 0.0672 0.0002 0.0005 0.0017
Communities and Crime ($n=101$) 0.0059 0.0123 0.0282 0.0484 0.0547 0.0602 0.0005 0.0008 0.0014
TUANDROMD ($n=241$) 0.0050 0.0155 0.0431 0.0515 0.0553 0.0612 0.0002 0.0012 0.0022
Table 1: The relative squared $\ell_{2}$-norm error (i.e., $\|\tilde{\boldsymbol{\phi}}-\boldsymbol{\phi}\|_{2}^{2}/\|\boldsymbol{\phi}\|_{2}^{2}$) of Banzhaf estimators with $m=20n$ samples over 50 runs. We bold the lowest error for the 25% percentile, median, and 75% percentile. Across all eight datasets and every summary statistic, Kernel Banzhaf returns estimates with the lowest error, typically by an order of magnitude.
Time Complexity.
Let $T$ be the time complexity of evaluating the set function $v$. The total runtime of our Kernel Banzhaf method is $O(Tm+mn^{2})$. In particular, it takes $O(n)$ time to sample a random set and its compliment, for a total of $O(nm)$ cost to construct $\tilde{\mathbf{A}}$. It then takes $O(Tm)$ time to construct $\tilde{\mathbf{b}}$, and finally $O(mn^{2})$ time to solve the $m\times n$ least squared problem involving $\tilde{\mathbf{A}}$ and $\tilde{\mathbf{b}}$. For most applications in ML and explainable AI, we expect the $O(Tm)$ term to dominate. For example, even a forward pass on a shallow fully connected neural network with $n$ features takes $O(n^{2})$ time, so evaluating any attribution-based value function takes at least $O(mn^{2})$ time in total. Figure 7 in Appendix F confirms that the cost of Kernel Banzhaf (and prior Monte Carlo methods) is dominated by the number of samples, $m$, i.e., the number of evaluations of $v$.
@@ 3.3 Approximation Guarantees
Despite its relative simplicity, it is possible to prove strong theoretically guarantees on the approximation quality of the Kernel Banzaf method. Such guarantees have recently been show for regression-based Shapley evaluation using results on subsampled regression problems from randomized numerical linear algebra (Musco & Witter, 2024; Woodruff et al., 2014; Drineas & Mahoney, 2018). In Appendix A, we use similar tools to first prove that, using a near-linear number of samples, we can solve the Banzhaf value regression formulation from Theorem 3.1 near optimally:
@@
If $m=O(n\log\frac{n}{\delta}+\frac{n}{\delta\epsilon})$, Algorithm 1 produces an estimate $\tilde{\boldsymbol{\phi}}$ that satisfies, with probability $1-\delta$,
$\|\mathbf{A}\tilde{\boldsymbol{\phi}}-\mathbf{b}\|_{2}^{2}\leq(1+\epsilon)\|\mathbf{A}\boldsymbol{\phi}-\mathbf{b}\|_{2}^{2}.$ (7)
Due to a known lower bound on the sample complexity of agnostic active linear regression (Chen & Price, 2019), Theorem 3.2 can only be improved in the logarithmic factor on $n$ (Chen & Price, 2019). For constant $\epsilon$ and $\delta$ this is intuitive: in particular, consider a linear set function where $\mathbf{b}$ is in the span of $\mathbf{A}$. Then the right-hand side is 0 and we must recover $\boldsymbol{\phi}$ exactly. To do this, we need to observe at least $n$ linearly independent rows of the regression problem.
As a direct consequence of Theorem 3.2, we obtain a bound on the Euclidean distances between the approximate Banzhaf values computed by our Kernel Banzhaf method and the true values. In particular, in Appendix A, we show:
Corollary 3.3.
Let $\gamma=\|\mathbf{A}\boldsymbol{\phi}-\mathbf{b}\|_{2}^{2}/\|\mathbf{A}\boldsymbol{\phi}\|_{2}^{2}$. Any $\tilde{\boldsymbol{\phi}}$ that satisfies Equation 7 also satisfies
$\|\tilde{\boldsymbol{\phi}}-\boldsymbol{\phi}\|_{2}^{2}\leq\epsilon\gamma\|\boldsymbol{\phi}\|_{2}^{2}.$ (8)
The reverse is also true: Equation 8 implies Equation 7.
Figure 2: Plots comparing the relative squared $\ell_{2}$-norm error (i.e., $\|\tilde{\boldsymbol{\phi}}-\boldsymbol{\phi}\|_{2}^{2}/\|\boldsymbol{\phi}\|_{2}^{2}$) for each Banzhaf estimator across increasing sample sizes in eight datasets. Each point represents the median of 50 runs, with shaded areas indicating the 25th to 75th percentiles.
Corollary 3.3 incurs a dependence on a parameter $\gamma$, which can be viewed as capturing how close $v$ is to an additive set function (i.e., how close $\mathbf{b}$ is to a linear function in the rows of $\mathbf{A}$, which can be viewed as set indicator vectors). Since there is no relaxation in the guarantee from Theorem 3.2 to Corollary 3.3 and, as discussed, the sample complexity of Theorem 3.2 is near optimal, we believe a dependence on $\gamma$ is inherent to the Kernal Banzhaf method. Fortunately, we find experimentally that the quantity is typically small – the median value of $\gamma$ was around $100$ in our experiments on datasets that were small enough to compute the quantity. In an idea case, we could even have $\gamma=0$ when $\mathbf{b}$ lies in the span of $\mathbf{A}$, and we would recover the exact Banzhaf values.
Comparison of Theoretical Guarantees.
Basic concentration bounds can be used to prove similar guarantees for the Monte Carlo and Maximum Sample Reuse estimators (see e.g., Theorems 4.8 and 4.9 in Wang & Jia (2023)). When MC is run with $O\left(\frac{n^{2}}{\epsilon}\log(\frac{n}{\delta})\right)$ samples and when MSR is run with $O\left(\frac{n}{\epsilon}\log(\frac{n}{\delta})\right)$ samples, they respectively return estimates $\tilde{\boldsymbol{\phi}}$ that satisfy, with high probability,
$\|\tilde{\boldsymbol{\phi}}-\boldsymbol{\phi}\|_{2}^{2}\leq\epsilon\max_{S\subseteq[n]}v(S)^{2}.$
That is, both prior methods obtain a bound that depends on the maximum of $v$, rather than $\|\boldsymbol{\phi}\|_{2}^{2}$. In general, we expect the maximum magnitude of the set function to be larger than the Banzhaf values, which measure an average marginal contribution. This is reflected in the superior Kernel Banzhaf in our experiments (see e.g., Table 1).
3.4 Extension to Probabilistic Values
Banzhaf and Shapley values can be generalized to so-called probabilistic values, which have also found applications in machine learning (Kwon & Zou, 2022; Li & Yu, 2024a, b). In this section, we show how our regression formulation can be extended to probabilistic values.
For a weight vector $\mathbf{p}\in[0,1]^{n}$ such that $\sum_{\ell=0}^{n-1}\binom{n-1}{\ell}p_{\ell}=1$, the probabilistic value is given by
$\phi^{\textnormal{prob}}_{i}=\sum_{S\subseteq[n]\setminus\{i\}}p_{|S|}[v(S\cup\{i\})-v(S)].$
Define the quantities $a_{n}$ and $b_{n}$ as follows:
$a_{n}$ :$\vcentcolon=2\sum_{\ell=0}^{n-1}\binom{n-1}{\ell}p_{\ell}^{2}-\sum_{\ell=1}^{n-1}\binom{n-2}{\ell-1}\left(p_{\ell}-p_{\ell-1}\right)^{2}$ (9)
$b_{n}$ :$\vcentcolon=\frac{1}{a_{n}}\sum_{\ell=1}^{n-1}\binom{n-2}{\ell-1}\left(p_{\ell}-p_{\ell-1}\right)^{2}.$ (10)
Note that both $a_{n}$ and $b_{n}$ depend only on $n$ and $\mathbf{p}$, and can be computed in $O(n)$ time. In the Banzhaf setting when $p_{\ell}=p_{\ell-1}$ for all $\ell$, observe that $b_{n}=0$.
Construct a design matrix be $\mathbf{A}^{\textnormal{prob}}\in\mathbb{R}^{2^{n}\times n}$ with
$[\mathbf{A}^{\textnormal{prob}}]_{S,i}=\begin{cases}\frac{p_{|S|-1}}{a_{n}}&\textrm{if }i\in S\\ -\frac{p_{|S|}}{a_{n}}&\textrm{if }i\notin S.\end{cases}$
Exactly as we did for Banzhaf values, let $\mathbf{b}^{\textnormal{prob}}_{S}=v(S)$.
Lemma 3.4 (Extended Regression Equivalence).
Let
$\mathbf{x}^{*}=\operatorname*{arg\,min}_{\mathbf{x}\in\mathbb{R}^{n}}\|\mathbf{A}^{\textnormal{prob}}\mathbf{x}-\mathbf{b}^{\textnormal{prob}}\|_{2}^{2}.$
Then $\boldsymbol{\phi}^{\textnormal{prob}}=(\mathbf{I}+b_{n}\mathbf{11^{\top}})\mathbf{x}^{*}$.
We defer the proof of Lemma 3.4 to Appendix B.
Like in the Banzhaf setting, we can sample the full regression problem to produce a subsampled design matrix $\tilde{\mathbf{A}}^{\textnormal{prob}}\in\mathbb{R}^{m\times n}$ and target vector $\tilde{\mathbf{b}}^{\textnormal{prob}}\in\mathbb{R}^{m}$. Let ${\tilde{\boldsymbol{\phi}}}^{\textnormal{prob}}=\operatorname*{arg\,min}_{\mathbf{x}}\|\tilde{\mathbf{A}}^{\textnormal{prob}}\mathbf{x}-\tilde{\mathbf{b}}^{\textnormal{prob}}\|_{2}^{2}$ be the solution to this subsampled problem. The following theorem (proven in Appendix B) gives guarantees on the accuracy of this solution.
Figure 3: Plots of relative squared $\ell_{2}$-norm error by noise level across Banzhaf estimators. For each noise level $\sigma$, the estimator observes $v(S)+x$ where $x\sim\mathcal{N}(0,\sigma)$. Kernel Banzhaf outperforms for lower noise levels, eventually matching its ablated version and MSR for larger noise. MC is worse for all noise settings, likely because its constituent estimates are $v(S\cup\{i\})+x-v(S)-x^{\prime}$, increasing the variance of the noise.
Theorem 3.5 (Probabilistic Value Approximation).
Let $\gamma=\|\mathbf{A}^{\textnormal{prob}}\mathbf{x}^{*}-\mathbf{b}^{\textnormal{prob}}\|_{2}^{2}/\|\mathbf{A}^{\textnormal{prob}}\mathbf{x}^{*}\|_{2}^{2}$. There is a sampling method that takes $m=O(n\log(n/\delta)+n\log(1/\delta)/\epsilon)$ samples and ensures that, with probability $1-\delta$,
$\|\boldsymbol{\phi}^{\textnormal{prob}}-{\tilde{\boldsymbol{\phi}}}^{\textnormal{prob}}\|_{2}^{2}\leq\epsilon\gamma(1+nb_{n})^{3}\|\boldsymbol{\phi}^{\textnormal{prob}}\|_{2}^{2}.$
Note that Theorem 3.5 recovers Corollary 3.3 for Banzhaf values since $b_{n}=0$. Initial experiments suggest that the dependence on $b_{n}$ can limit the accuracy of this regression formualation for other probabilistic values, since we could have $b_{n}\gg 0$. However, we leave further exploration of the topic to future work.
4 Experiments
We conduct extensive experimental evaluations of Banzhaf estimators across eight popular datasets.  [^1] Our code is available at https://github.com/lyrain2001/kernel-banzhaf. In these experiments, we focus on feature attribution tasks for two reasons: First, we are interested in explainable AI applications where users seek to interpret model predictions in terms of their input features. Second, we can efficiently compute Banzhaf values even for large $n$, provided the set function corresponds to the feature attribution function of a decision tree model (Karczmarz et al., 2022). This allows us to obtain a baseline, and thus an accurate calculation of the accuracy of various Banzhaf values estimation methods.
In addition to experiments estimating the Banzhaf values of tree-based models, we also explore neural network models using smaller datasets for which Equation 2 can be used to calculate exact Banzhaf values. These neural network experiments are presented in Figure 5 of Appendix D, and align with our findings for decision trees.
We measure approximation error with $\|\tilde{\boldsymbol{\phi}}-\boldsymbol{\phi}\|_{2}^{2}/\|\boldsymbol{\phi}\|_{2}^{2}$ i.e., the relative squared $\ell_{2}$-norm error between the estimated and true Banzhaf values. In Appendix E, we also evaluate estimators in terms of the objective value $\|\mathbf{A\tilde{\boldsymbol{\phi}}-b}\|_{2}$ suggested by the linear regression formulation; the results in Figure 6 corroborate our findings on the $\ell_{2}$-norm error.
We provide more details on the datasets and implementation details in Appendix C.
Figure 4: Comparison of top 20% feature ranking recovery using Cayley distance (lower is more accurate). Recovering the most impactful features is an primary component of comparing and selecting the most important features for machine learnign models. Especially as the sizes of the datasets increase, Kernel Banzhaf gives the best performance.
4.1 Baselines
We compare Kernel Banzhaf with the following methods:
MC The Monte Carlo (MC) algorithm estimates each Banzhaf value individually. Let $\mathcal{S}_{i}$ be the subsets sampled for the estimate of the $i$th Banzhaf value. We have that $\sum_{i=1}^{m}|\mathcal{S}_{i}|=m$. The MC estimate for player $i$ is given by
$\tilde{\phi}^{\textrm{MC}}_{i}=\frac{1}{|\mathcal{S}_{i}|}\sum_{S\in\mathcal{S}_{i}}[v(S\cup\{i\})-v(S)].$
$\tilde{\phi}^{\textrm{MSR}}_{i}=\frac{1}{|\mathcal{S}_{\ni i}|}\sum_{S\in\mathcal{S}_{\ni i}}v(S)-\frac{1}{|\mathcal{S}_{\not\owns i}|}\sum_{S\in\mathcal{S}_{\not\owns i}}v(S).$
Kernel Banzhaf (Excluding Pairs) This algorithm is Kernel Banzhaf (Algorithm 1) but without paired sampling.
4.2 Sample Efficiency
Since evaluating the set function $v$ generally dominates the estimator’s time complexity (see confirmation in Appendix F), we investigate estimator error by number of samples in Figure 2. As expected, the error for all estimators decreases as the sample size increases. In terms of performance, Kernel Banzhaf gives the lowest error, followed by its ablated variant without paired sampling, and then MC. In Wang & Jia (2023), the MSR algorithm was shown to converge faster than the MC algorithm, but faster convergence does not always imply better accuracy.
4.3 Robustness to Noise
The set functions in explainable AI tasks are often estimates of a random process e.g., the expectation of a model’s prediction on a randomly sampled observation. We measure how robust each estimator is to inaccurate set function values by running experiments on the modified set function $v(S)+x$ where $x\sim\mathcal{N}(0,\sigma)$ for a given noise level $\sigma$. As shown in Figure 3, Kernel Banzhaf and its variant without paired sampling are more robust to this noise than the other algorithms, especially when the level of noise size is low. This suggests that Kernel Banzhaf is particularly well-suited for real-world applications where the set function is approximated. In Appendix I, we explore how adversarial noise affects estimator performance, finding that Kernel Banzhaf continues to give the best performance in the adversarial noise setting.
4.4 Feature Ranking Recovery
Aside from evaluating the quantitative errors between estimated and exact Banzhaf values, we also evaluate how well each estimator recovers the rank of the most important features. Such ranking tasks, where we are interested in recovering the rank rather than quantitative value of a feature, is a principal component of comparing and selecting important features for machine learning models. Figure 4 shows the Cayley distance between the estimated ranks of the top 20% of features and the true ranks. Especially for datasets with larger $n$, Kernel Banzhaf gives the most accurate rankings. We corroborate our findings in terms of the Spearman correlation coefficient (where a higher value is better) in Figure 9 of Appendix H.
5 Conclusion
In this work, we present Kernel Banzhaf, the first regression-based estimator for Banzhaf values. Our experiments suggest that Kernel Banzhaf significantly outperforms existing methods for estimating Banzhaf values. Our theoretical analysis establishes robust approximation guarantees that depend on the size of the Banzhaf values, in contrast to Monte Carlo estimators which depend on the magnitude of the underlying set function.
One avenue for future work is to evaluate the generalization of Kernel Banzhaf on other probabilistic values. However, it is not clear how to efficiently compute general probabilistic values and the poor approximation guarantee when $b_{n}\neq 0$ may be reflected in the generalized algorithm’s performance.
Impact Statement
This paper presents a new algorithm for more efficiently computing scores that are widely used to better understand decisions made by black-box machine learning algorithms. While all possible applications of the work are difficult to predict, we thus expect the results to have a net positive impact on the ethical and responsible application of machine learning in society.
Acknowledgements
Freire and Liu were partially supported by NSF Awards IIS-2106888 and the DARPA ASKEM and ARPA-H BDF programs. Liu initiated this project during her summer internship at Google Research. Witter was supported by the National Science Foundation under Graduate Research Fellowship Grant No. DGE-2234660. Musco was partially supported by NSF Award CCF-2045590. Opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not reflect the views of NSF or DARPA.
Appendix A Kernel Banzhaf Approximation Guarantees
In this section, we prove theoretical guarantees on the performance of Kernel Banzhaf. We begin with a standard regression sampling result, the proof of which we defer to the end of this section.
Theorem A.1.
Consider a subsampling matrix $\mathbf{S}\in\mathbb{R}^{m\times 2^{n}}$ where each row contains exactly one non-zero entry, selected (possibly in a paired way) with probability proportional to the leverage scores of the corresponding row in $\mathbf{A}$. Define $\tilde{\mathbf{A}}=\mathbf{SA}$ and $\tilde{\mathbf{b}}=\mathbf{Sb}$. Let the solution to the subsampled problem be $\tilde{\boldsymbol{\phi}}=\operatorname*{arg\,min}_{\mathbf{x}}\|\tilde{\mathbf{A}}\mathbf{x}-\tilde{\mathbf{b}}\|_{2}^{2}$. If $m=O(n\log(\frac{n}{\delta})+\frac{n}{\delta\epsilon})$, then
$\|\mathbf{A}\tilde{\boldsymbol{\phi}}-\mathbf{b}\|_{2}^{2}\leq(1+\epsilon)\|\mathbf{A}\boldsymbol{\phi}-\mathbf{b}\|_{2}^{2}$ (11)